Three lottery tickets for first, second, and third prizes are drawn from a group of 45 tickets

1
Expert's answer
2022-06-23T18:07:21-0400

The Number of sample points for awarding the three process is the product of the number of possible outcomes after each pick.

Total number of tickets = 45

First draw: total possible outcomes = 45

Second draw: total possible outcomes = (45 - 1) = 44

Third draw: total possible outcomes = (44 - 1)=43

The total number of sample points can be calculated thus:


"45(44)(43)=85140"


Therefore, the total number of sample points is 85140.